Your teacher is giving you a test worth 100 points and containing 40 questions. There are 2-point and 4-point questions on the test. How many type of question are on the test?
Answer by ankor@dixie-net.com(12692)

Your teacher is giving you a test worth 100 points and containing 40 questions. There are 2-point and 4-point questions on the test. How many type of question are on the test?
:
Let x = number of 4 pt questions; Then (40-x) = number of 2 pt questions;
:
A simple equation;
4x + 2(40-x) = 100
4x + 80 - 2x = 100
4x - 2x = 100 - 80
2x = 20
x = 20/2 = 10 4 pt questions
:
40 - 10 - 30 2 pt questions
:
Check 4(10) + 2(30) = 100